# Hedgehog topological defects in 3D amorphous solids

### v1

@@ Abstract
The underlying structural disorder renders the concept of topological defects in amorphous solids difficult to apply and hinders a first-principle identification of the microscopic carriers of plasticity and of the regions more prone to structural rearrangements (“soft spots”). Recently, it has been proposed that well-defined topological defects can still be identified in glasses, and correlated to local and global plasticity, by looking at the eigenvector field or the particle displacement field. Nevertheless, all the existing proposals and analyses are only valid in two spatial dimensions. In this work, we propose the idea of using hedgehog topological defects to characterize the plasticity of 3D glasses and to geometrically predict the location of their soft spots. We corroborate our proposal by simulating a Kremer-Grest 3D polymer glass, and by using both the normal mode eigenvector field and the displacement field around large plastic events. Our results confirm that a topological characterization of plasticity in glasses is feasible, and provide a concrete realization of this program in 3D amorphous systems.
@@
Introduction – Topological defects, such as dislocations and disclinations, play a pivotal role for the description of mechanical failure in crystalline solids since they provide the microscopic origin of their plasticity Taylor (1934); Polanyi (1934); Orowan (1934). Topological defects also provide the building blocks for the theory of crystal melting in two dimensions Kosterlitz and Thouless (1973); Halperin and Nelson (1978); Young (1979). Despite several attempts Spaepen (1978); Cotterill (1979); Blackett (1979); Morris (1979); Popescu (1984); Egami and Vitek (1984); Shi (1993); Chaudhari et al. (1979); Egami et al. (1980); Acharya and Widom (2017); Nelson (1983); Steinhardt and Chaudhari (1981), topological defects associated to translational symmetry (dislocations and disclinations) cannot be formally defined within the structure of amorphous solids such as glasses, because of the absence of translational long-range order. This difficulty remains the main cause for the lack of a microscopic theory of plasticity and mechanical deformations in glasses Hufnagel et al. (2016); Rodney et al. (2011); Zaccone (2023).
@@
More in general, the definition and identification of “defects” constitute the main obstacle for establishing a robust causal relation between structure and dynamics in amorphous solids (see e.g. Wang (2019); Cheng and Ma (2011)). So far, most of the approaches to solve this problem are still based on ad-hoc coarse-grained descriptions Nicolas et al. (2018), vibrational soft modes Rottler et al. (2014); Manning and Liu (2011); Ding et al. (2014); Chen et al. (2011); Kapteijns et al. (2020); Tanguy et al. (2010); Smessaert and Rottler (2014a), shear transformation zones revisited under the lens of anharmonic quasi-localized modes Zylberg et al. (2017); Richard et al. (2021); Gartner and Lerner (2016); Ruan et al. (2022); Xu et al. (2018, 2021), phenomenological structural indicators Richard et al. (2020); Widmer-Cooper et al. (2008) or machine-learning algorithms Fan and Ma (2021); Ciarella et al. (2023).
@@
Two recent works Baggioli et al. (2021); Wu et al. (2023) (see also Baggioli (2023)) have shown that not only topological concepts can be applied (albeit with proper generalizations) to amorphous solids, as done for vortices in superfluids or dislocations in crystals, but that they can be also extremely useful to predict their mechanical properties both on a local and a global scale.
@@
A first approach Baggioli et al. (2022) is based on the idea of continuous Burger vector Kleman and Friedel (2008) applied to the displacement field, and has been shown to be able to accurately locate plastic events under deformation and the yielding point Baggioli et al. (2021). It is likely that this concept is related to the geometric charges previously discussed as the carriers of plastic screening in glasses Lemaître et al. (2021). The second approach Wu et al. (2023) relies on the identification of vortex-like defects in the eigenvector field that strongly correlate with the plastic spots and display tight connections with the widely used concept of shear transformation zones Desmarchelier et al. (2024). This second type of vortex-like topological defects have been recently experimentally identified in a 2D colloidal glass Vaibhav et al. (2024), confirming the simulation results.
@@
Importantly, all the proposals and analyses so far apply only to two-dimensional (2D) amorphous solids and lack direct application to more realistic 3D scenarios. In Cao et al. (2018), disclination-like defects, defined using polytetrahedral order Nelson and Spaepen (1989), were considered in a 3D experimental granular matter system and were shown to correlate with plasticity. In Bera et al. (2024), a 2D slicing method has been used to study vortex-like topological defects in a 3D simulation polymer glass revealing a strong clustering tendency of negative charged topological defects right before large plastic events.
@@
Hedgehog topological defects – Given the large variety of topological defects that can be defined within a three-dimensional vector field (point defects, line defects and surface defects) Kleinert (1989); Nelson (2002), we suggest that hedgehog topological defects (HTD) are the simplest choice to start with. We adapt the definition of HTDs from nematic liquid crystals Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023) and Heisenberg magnets Fumeron and Berche (2023); Berg and Lüscher (1989); Holm and Janke (1994); Berg and Lüscher (1981) to the eigenvector and displacement fields of a 3D simulation glass.
For a generic 3D continuous vector field $\vec{s}$, we define the hedgehog topological charge as Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023)
[…]
where $\Sigma$ is an arbitrary closed 2D surface parametrized by coordinates $(x_{1},x_{2})$. $Q$ represents a topological invariant in the sense that it does not change under continuous and smooth deformations of the manifold.
Figure 1: (a) Discrete vector at a simple cubic lattice corners illustrating a radially outward field for a monopole with $Q=+1$ (red). A vertex $x^{*}$ is shown where $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ field at four corners of the vertex. (b) Continuous vector field (left) and discrete vector field (right) for a monopole. (c) Radially inward field representing an anti-monopole with $Q=-1$ (blue). (d) Continuous (left) and discrete (right) vector fields for $Q=-1$.
From our simulations, we first obtain a discrete vector field ${\vec{s}}_{l,m,n}$ at each lattice grid point $(l,m,n)$ in a simple cubic lattice. More details are provided in the Supplementary Material (SM). Let $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$, and $\vec{s}_{4}$ be the unit vectors of the field associated with one face of the cube, where $\vec{s}_{1}$, $\vec{s}_{2}$, and $\vec{s}_{3}$ are positioned at the corners of triangle $i$ on that face, chosen such that the sequence $1-2-3-1$ corresponds to a counterclockwise rotation along the outward normal to the surface of the triangle. See Fig. 1(a) for a representation of these vectors on the example of a radially outgoing field (monopole). The topological charge $Q$ enclosed by the unit cube is then expressed in a discrete fashion as Holm and Janke (1994); Berg and Lüscher (1981),
@@
where the sum is carried over all $6$ vertices of the unit cube and $x^{*}$ is the vertex associated with the four vectors $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ (as shown in Fig. 1(a)). Here, $q(x^{*})$ is expressed as
@@
where $(\sigma_{s}A)(\vec{s}_{1},\vec{s}_{2},\vec{s}_{3})$ is the signed area of the spherical (non-Euclidean) triangles with corners $\vec{s}_{1}$, $\vec{s}_{2}$ and $\vec{s}_{3}$. The sign of the area can be calculated as ${\sigma_{s}=\rm{sign}}\{\vec{s}_{1}\cdot[\vec{s}_{2}\times\vec{s}_{3}]\}$. Finally, the value of $A$ is calculated as Holm and Janke (1994); Berg and Lüscher (1981),
@@
A visual representation of this algorithm for the case of hedgehog defects with charges $Q=\pm 1$ (monopole and anti-monopole) is presented in Fig. 1.
3D polymer glass model – We simulate a 3D polymer glass system using the Kremer-Grest model Kremer and Grest (1990) with $N=10{,}000$ monomers of alternating masses of $m_{1}=1$ and $m_{2}=3$ Kriuchevskyi et al. (2020) interacting via the Lennard-Jones (LJ) potential ($u_{\text{LJ}}$) and the finite extensible nonlinear elastic (FENE) potential ($u_{\text{FENE}}$). We set the dihedral bending rigidity of the bonds to zero, as appropriate for flexible chains Palyulin et al. (2018). More details can be found in the SM.
@@
The eigenvalues $\lambda_{k}$ (or equivalently eigenfrequencies $\omega_{k}=\sqrt{\lambda_{k}}$) and their associated eigenvectors are obtained by direct diagonalization of the Hessian matrix. In the rest of the manuscript, we will use the symbols $\omega_{k}$ and $\omega$ interchangeably.
Soft spots from the eigenvector field – We start by applying our algorithm to the 3D eigenvector field of the simulated polymer glass system. To determine the HTDs, we derived the smoothed eigenvector field on a simple cubic lattice (see SM for details). For direct visualization, we present a snapshot of the smooth eigenvector field that exhibits an hedgehog anti-monopole and a monopole in the SM, Fig. S2.
By scanning the eigenvector field at different frequencies, we have verified that the total number of HTDs scales at low frequency (up to $\omega\approx 4$) as $N_{\text{HTD}}\propto\omega^{2}$ (shown in SM, Fig. S3(b)) and correlates with the corresponding density of states, which exhibits Debye scaling, $D(\omega)\propto\omega^{2}$ (see SM, Fig. S3(a)) at low frequency. This result is consistent with previous findings in 2D simulation and experimental works Wu et al. (2023); Vaibhav et al. (2024).
After confirming the validity of the 3D algorithm presented above, we investigate the physical properties of the HTDs, and in particular whether they bear any connection with plastic soft spots. In order to do so, we follow Smessaert and Rottler (2014b) and we identify the soft spots in the 3D model glass system using the softness parameter
@@
Figure 2: (a) The color map visualizes the 3D average topological charge density. The black symbols are the locations of soft spots, identified using the softness parameter, Eq. (6). (b)-(f) Pair correlation functions between HTDs: monopoles (p) and anti-monopoles (n), and plastic soft spots (s) for different frequencies. Note that the curves for $\omega>1.1$ are shifted upward by constant factors for better visualization in this and all other similar data sets. See SM, Fig. S4, for extended data.
@@
In Fig. 2(a), we show a 3D color map indicating the intensity of the average topological charge density (see SM for details). Red and blue colored regions represent therefore areas with higher concentrations of positive and negative topological charge $Q$, respectively. In the same box, we provide the locations of the plastic soft spots using black symbols. In order to test the correlation between HTDs and plasticity, in Fig. 2(b)-(g), we present various pair correlation functions between HTDs (p), anti-monopole HTDs (n) and plastic soft spots (s) for different frequencies.
@@
In Fig. 2(b)-(c), the correlation between HTDs and soft spots is shown. Our results reveal that both positive and negative defects exhibit strong correlations with soft spots at short and intermediate distances at low frequency, that nevertheless vanish quickly by increasing $\omega$ (see Fig. S4 in the SM). This outcome presents marked differences with the 2D case Wu et al. (2023), where soft spots predominantly correlate with negative defects at short distances. This might be attributed to the different nature of point defects in 2D and 3D systems Kleman and Lavrentovich (2003).
@@
Moreover, HTDs with same charge also display a peak in their spatial correlation at approximately the same distance, see Fig. 2(d) and Fig. 2(f), indicating a widespread tendency to clustering of HTDs. This may be explained with the fact that HTDs tend to nucleate more in regions of significant nonaffine motions Baggioli et al. (2022), which are spatially localized DiDonna and Lubensky (2005). Hence, it is statistically more likely to have charges, even with same sign, next to each other. Finally, in Fig. 2(e), we observe a very strong short-distance correlation between monopole and anti-monopole defects that suggests that defects with opposite charges attract each other and tend to pair together to minimize the local topological charge. This correlation is very strong at low frequency and becomes weaker at higher $\omega$.
@@
Predicting plasticity from the displacement vector – Inspired by the proposal of looking at topological defects in the dynamical displacement field $\vec{u}$ Baggioli et al. (2021), rather than in the eigenvector field, we now apply the same 3D algorithm to find HTDs in $\vec{u}$. We notice that there is no direct relation between the displacement field and a single eigenvector $\mathbf{e}_{k}$, despite $\vec{u}$ can be decomposed onto the complete basis set of eigenvectors. Therefore, a correlation between topological defects in the eigenvector field and in the displacement field is not obvious. Nevertheless, around plastic events, it is known that the dynamical response is dominated by a few localized and low frequency modes (see e.g. Yang et al. (2020)), providing a possible connection between the topology of the low-frequency eigenvectors and that of the displacement field.
@@
In our study, we apply quasi-static shear to our system using the athermal quasi static (AQS) protocol Lemaître and Maloney (2006) with strain step $\delta\gamma=10^{-3}$. The stress-strain curve is presented in the SM (Fig. S1(c)) and displays a yielding instability at $\gamma^{*}\approx 0.099$. For presentation purposes, here we focus on four plastic events, including the yielding point, and consider the non-affine displacement field at these specific strain values. In order to test the universality of our findings, we define the soft spots using two different methods: (I) the softness parameter defined in Eq. (6) (and used in the previous analysis) and (II) the value of the particle level non-affine displacement field $\vec{u}_{\text{NA}}$. Within this second method, we identify the plastic soft spots by looking at the locations of the $5\%$ of particles having higher values of $|\vec{u}_{\text{NA}}|$.
@@
Figure 3: (a)-(d) Spatial correlation functions between soft spots (identified using the softness parameter, Eq. (6)) and HTDs within the non-affine displacement field across four plastic events at different $\gamma$. Correlations between soft spots and HTDs with $Q=+1$($-1$) are shown in red (blue).

### v2

@@ Abstract
The underlying structural disorder renders the concept of topological defects in amorphous solids difficult to apply and hinders a first-principle identification of the microscopic carriers of plasticity and of the regions more prone to structural rearrangements (“soft spots”). Recently, it has been proposed that well-defined topological defects can still be identified in glasses, and correlated to local and global plasticity, by looking at the eigenvector field or the particle displacement field. Nevertheless, all the existing proposals and analyses are only valid in two spatial dimensions. In this work, we propose the idea of using hedgehog topological defects to characterize the plasticity of 3D glasses and to geometrically predict the location of their soft spots. We corroborate our proposal by simulating a Kremer-Grest 3D polymer glass, and by using both the normal mode eigenvector field and the displacement field around large plastic events. Contrary to the 2D case, the sign of the topological charge defined from the eigenvector field is ambiguous and the geometry of the topological defects, whether radial or hyperbolic, plays a fundamental role in 3D. In fact, we find that the topological hedgehog defects relevant for plasticity are those exhibiting hyperbolic geometry, resembling the saddle-point structure of 2D topological defects with negative winding number (anti-vortices). Our results confirm that a topological characterization of plasticity in glasses is feasible, and provide a concrete realization of this program in 3D amorphous systems.
@@
Topological defects, such as dislocations and disclinations, play a pivotal role for the description of mechanical failure in crystalline solids since they provide the microscopic origin of their plasticity [1, 2, 3]. Topological defects also provide the building blocks for the theory of crystal melting in two dimensions [4, 5, 6]. Despite several attempts [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], topological defects associated to translational symmetry (e.g., dislocations) cannot be formally defined within the structure of amorphous solids such as glasses, because of the absence of translational long-range order (or quasi-long range order in 2D systems). This difficulty remains the main cause for the lack of a microscopic theory of plasticity and mechanical deformations in glasses [19, 20, 21].
@@
More in general, the definition and identification of “defects” constitute the main obstacle for establishing a robust causal relation between structure and dynamics in amorphous solids (see e.g. [22, 23]). So far, most of the approaches to solve this problem are still based on ad-hoc coarse-grained descriptions [24], vibrational soft modes [25, 26, 27, 28, 29, 30, 31], shear transformation zones revisited under the lens of anharmonic quasi-localized modes [32, 33, 34, 35, 36, 37], phenomenological structural indicators [38, 39] or machine-learning algorithms [40, 41, 42, 43].
@@
Two recent works [44, 45] (see also [46]) have shown that not only topological concepts can be applied (albeit with proper generalizations) to amorphous solids, as done for vortices in superfluids or dislocations in crystals, but that they can be also extremely useful to predict their mechanical properties both on a local and a global scale.
@@
A first approach [47] is based on the idea of continuous Burgers vector [48] applied to the displacement field, and has been shown to be able to accurately locate plastic events under deformation and the yielding point [44]. It is likely that this concept is related to the geometric charges previously discussed as the carriers of plastic screening in glasses [49]. The second approach [45] relies on the identification of vortex-like defects in the eigenvector field that strongly correlate with the plastic spots and display tight connections with the widely used concept of shear transformation zones [50]. This second type of vortex-like topological defects have been recently experimentally identified in a 2D colloidal glass [51], confirming the simulation results.
@@
Importantly, all the proposals and analyses so far apply only to two-dimensional (2D) amorphous solids and lack direct application to more realistic 3D scenarios. In [52], disclination-like defects, defined using polytetrahedral order [53], were considered in a 3D experimental granular matter system and were shown to correlate with plasticity. In [54], a 2D slicing method has been used to study vortex-like topological defects in a 3D simulation polymer glass revealing a strong clustering tendency of negative charged topological defects right before large plastic events.
@@
Hedgehog Topological Defects and 3D polymer glass model
Given the large variety of topological defects that can be defined within a three-dimensional vector field (point defects, line defects and surface defects) [55, 56], we suggest that hedgehog topological defects (HTD) are the simplest choice to start with. We adapt the definition of HTDs from nematic liquid crystals [57, 58, 59] and Heisenberg magnets [60, 61, 62, 63] to the eigenvector and displacement fields of a 3D simulation glass.
For a 3D unit vector field $\vec{s}(x)$ we define the hedgehog topological charge as [57, 58, 59, 64]
$Q=\frac{1}{4\pi}\int_{\Sigma}d^{2}x\leavevmode\nobreak\ \vec{s}\cdot(\partial_{1}\vec{s}\times\partial_{2}\vec{s}),$ (1)
where $\Sigma$ is an arbitrary closed 2D surface parametrized by coordinates $(x_{1},x_{2})$. $Q$ represents a topological invariant in the sense that it does not change under continuous and smooth deformations of the manifold. We notice that, whenever the system is invariant under the parity transformation $\vec{s}\rightarrow-\vec{s}$ (e.g., nematic liquid crystals), then the sign of the topological charge $Q$ is ambiguous and hence not physical.
Figure 1: Schematic representation of 3D point defects. a Discrete vector at a simple cubic lattice corners illustrating a radially outward field for a monopole with $Q=+1$ (red). A vertex $x^{*}$ is shown where $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ field at four corners of the vertex. b Continuous vector field (left) and discrete vector field (right) for a monopole. c Radially inward field representing an anti-monopole with $Q=-1$ (blue). d Continuous (left) and discrete (right) vector fields for $Q=-1$. e Radial defects with $Q=\pm 1$ are shown along with the 2D winding numbers of the projected vectors onto the faces of a cube. Negative 2D winding number ($q_{2D}=-1$) is indicated in orange, positive ($q_{2D}=+1$) in green. Stars mark the locations of the 2D point defects, while the corresponding surfaces are colored according to their winding number. f Hyperbolic defects with $Q=\pm 1$ are displayed with projected vectors and their associated winding numbers, using the same color scheme as in panel e.
We implement the concept of 3D point defect on a lattice following the methods in refs. [62, 63]. From our simulations, we first obtain the unit vector field ${\vec{s}}_{l,m,n}$ at each lattice grid point $(l,m,n)$ in a simple cubic lattice. More details are provided in the Methods section. Let $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$, and $\vec{s}_{4}$ be the unit vectors of the field associated with one face of the cube, where $\vec{s}_{1}$, $\vec{s}_{2}$, and $\vec{s}_{3}$ are positioned at the corners of triangle $i$ on that face, chosen such that the sequence $1-2-3-1$ corresponds to a counterclockwise rotation along the outward normal to the surface of the triangle. See Figure 1a for a representation of these vectors on the example of a radially outgoing field (monopole). The topological charge $Q$ enclosed by the unit cube is then expressed in a discrete fashion as [62, 63],
@@
where the sum is carried over all $6$ vertices of the unit cube and $x^{*}$ is the vertex associated with the four vectors $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ (as shown in Figure 1a). Here, $q(x^{*})$ is expressed as
@@
where $(\sigma_{s}A)(\vec{s}_{1},\vec{s}_{2},\vec{s}_{3})$ is the signed area of the spherical (non-Euclidean) triangles with corners $\vec{s}_{1}$, $\vec{s}_{2}$ and $\vec{s}_{3}$. The sign of the area can be calculated as ${\sigma_{s}=\rm{sign}}\{\vec{s}_{1}\cdot[\vec{s}_{2}\times\vec{s}_{3}]\}$. Finally, the value of $A$ is calculated as [62, 63],
@@
A visual representation of this algorithm for the case of hedgehog defects with charges $Q=\pm 1$ (monopole and anti-monopole) is presented in Figure 1a-d.
To identify the hedgehog defects, we use the cubic lattice construction as described above. We further analyze and classify the hedgehog defects based on the geometric structure of the vector field around their core [64, 65]. The vector fields for an “ideal” radial (R) defect (Figure 1a) and an “ideal” hyperbolic (H) defect (Figure 1b) are given by [65]:
$\vec{s}_{R}(x,y,z)=\{x,y,z\}(x^{2}+y^{2}+z^{2})^{-1/2},\qquad\vec{s}_{H}(x,y,z)=\{-x,-y,z\}(x^{2}+y^{2}+z^{2})^{-1/2},$ (5)
where both defects are located at the origin $(0,0,0)$. Importantly, both configurations in Eq.  (5) are associated to a topological charge $Q=+1$, proving that the sign of $Q$ does not fully characterize the defect properties in 3D since it is insensitive to the geometric structure. This is different to the case in 2D, where the sign of the winding number characterizes also the geometry of the defect: radial for vortices with positive charge, while hyperbolic for anti-vortices with negative charge. We notice that inverting the field direction in Eq. (5) ($\vec{s}\rightarrow-\vec{s}$) results in switching the sign of the topological charge $Q$ (as shown in Figure 1c,d). Nevertheless, this operation does not alter the radial or hyperbolic nature of the defect.
To practically characterize the geometric structure of the hedgehog defects, we project the 3D vector field onto the six faces of a cubic cell and then we compute the 2D winding number on each face as $q_{2D}=(1/2\pi)\oint\vec{\nabla}\theta\cdot\vec{d\ell}$, where $\theta$ is the phase angle and $\vec{d\ell}$ is the line element along a closed square loop. For an ideal radial hedgehog with $Q=+1$, then $q_{2D}=+1$ on all faces (Figure 1e). In contrast, an ideal hyperbolic defect with $Q=+1$ exhibits $q_{2D}=+1$ on two faces and $q_{2D}=-1$ on the remaining four (Figure 1f). This distinction allows us to clearly differentiate between radial and hyperbolic defects.
In our simulation system, the situation is more complicated. In fact, for an arbitrary defect core, we obtain various combinations of $q_{2D}$, which typically takes three distinct values, $-1$, 00 and $+1$ (for concrete examples, see Fig. 9 in the SM) and do not necessarily align with the ideal scenarios presented above. For each hedgehog defect, we further define the parameter $N_{s}$ that is the number of faces with $q_{2D}=-1$ (saddle-like). Clearly, for the ideal radial hedgehog $N_{s}=0$, and for the ideal hyperbolic hedgehog $N_{s}=4$. Then, we define the defects to be radial if $N_{s}=0$ and hyperbolic if $N_{s}>0$, while the value of $N_{s}$ could be thought as how exact the hyperbolic nature is. This classification is motivated by the inherent association of hyperbolic structures with saddle-like features onto the projected plane.
We simulate a 3D polymer glass system using the Kremer-Grest model [66] with $N=10{,}000$ monomers of alternating masses of $m_{1}=1$ and $m_{2}=3$ [67] interacting via the Lennard-Jones (LJ) potential ($u_{\text{LJ}}$) and the finite extensible nonlinear elastic (FENE) potential ($u_{\text{FENE}}$). We set the dihedral bending rigidity of the bonds to zero, as appropriate for flexible chains [68]. More details can be found in the Method section.
@@
The eigenvalues $\lambda_{k}$ (or equivalently the eigenfrequencies $\omega_{k}=\sqrt{\lambda_{k}}$) and their associated eigenvectors are obtained by direct diagonalization of the Hessian matrix. In the rest of the manuscript, we will use the symbols $\omega_{k}$ and $\omega$ interchangeably.
Soft spots from the topology of the eigenvector field
We start by applying our algorithm to the 3D eigenvector field of the simulated polymer glass system. To determine the HTDs, we derived the smoothed eigenvector field on a simple cubic lattice (see Methods for details). For direct visualization, we present a snapshot of the smooth eigenvector field that exhibits an hedgehog anti-monopole and a monopole in the Methods, Figure 6.
By scanning the eigenvector field at different frequencies, we have verified that the total number of HTDs scales at low frequency (up to $\omega\approx 4$) as $N_{\text{HTD}}\propto\omega^{2}$ (shown in Supplementary Material, Figure 8b) and correlates with the corresponding density of states, which exhibits Debye scaling, $D(\omega)\propto\omega^{2}$ (see SM, Figure 8b) at low frequency. We notice that this scaling does not align with a simple argument based on counting stationary points in plane wave eigenvectors presented in [45]. We found that such argument fails in our context since it reproduces only hedgehog defects with radial geometry, a scenario that appears incompatible with our simulation results. It would be interesting to understand this discrepancy better in the near future. Along these lines, we notice that a recent study [69] also reports a quadratic scaling of line defect density in 3D glasses.
After confirming the validity of the 3D algorithm presented above, we investigate the physical properties of the HTDs, and in particular whether they bear any connection with plastic soft spots. In order to do so, we follow [70] and we identify the soft spots in the 3D model glass system using the softness parameter
@@
Figure 2: Correlation between the soft spots and the topological defects. a The color map visualizes the 3D average topological charge density. The black symbols are the locations of soft spots, identified using the softness parameter, Eq. (7). b-f Pair correlation functions between HTDs: monopoles (p) and anti-monopoles (n), and plastic soft spots (s) for different frequencies. Note that the curves for $\omega>1.1$ are shifted upward by constant factors for better visualization in this and all other similar data sets. See SM, Fig. 9, for extended data.
Figure 3: Correlation between soft spots and topological defects with different geometric structure. a-d Pair correlation functions between HTDs with hyperbolic (H) and radial (R) nature and plastic soft spots (s) for different frequencies. The curves corresponding to $\omega>1.1$ are shifted upward by constant factors for improved visualization in this and all similar datasets. e The average correlation function between soft spots and hyperbolic (H) and radial (R) defects. f The average correlation function between soft spots and topological defects corresponding to different $N_{s}$ sectors. Error bars are included in all datasets.
@@
In Figure 2a, we show a 3D color map indicating the intensity of the average topological charge density (see SM for details). Red and blue colored regions represent therefore areas with higher concentrations of positive and negative topological charge $Q$, respectively. In the same box, we provide the locations of the plastic soft spots using black symbols. In order to test the correlation between HTDs and plasticity, in Figure 2b-g, we present various pair correlation functions between HTDs (p), anti-monopole HTDs (n) and plastic soft spots (s) for different frequencies.
@@
In Figure 2b-c, the correlation between HTDs and soft spots is shown. Our results reveal that both positive and negative defects exhibit strong correlations with soft spots at short and intermediate distances at low frequency, that nevertheless vanish quickly by increasing $\omega$ (see SM, Figure 9). This outcome presents marked differences with the 2D case [45], where soft spots predominantly correlate with negative defects at short distances. Here in 3D, we find that the sign of the topological charge $Q$ does not play an important role. In fact, this could have been anticipated by looking at the definition of the topological charge $Q$, Eq. (1), and by recalling that the sign of the eigenvector field in the normal mode analysis is arbitrary. Since inverting the sign of the eigenvector field is not physical, it follows that the sign of the topological charge $Q$, identifying the vector field $\vec{s}$ with the eigenvector field, is also not physical. This explains our findings that, up to statistical fluctuations, do not present any difference between positive and negative charges.
@@
Moreover, HTDs with same charge also display a peak in their spatial correlation at approximately the same distance, see Figure 2d and Figure 2f, indicating a widespread tendency to clustering of HTDs. This may be explained with the fact that HTDs tend to nucleate more in regions of significant nonaffine motions [47], which are spatially localized [71]. Hence, it is statistically more likely to have charges, even with same sign, next to each other. Finally, in Figure 2e, we observe a very strong short-distance correlation between monopole and anti-monopole defects that suggests that defects with opposite charges attract each other and tend to pair together to minimize the local topological charge. This correlation is very strong at low frequency and becomes weaker at higher $\omega$.
As discussed above, in 3D the sign of the charge $Q$ for point defects defined from the eigenvector field is arbitrary. Hence, differently from the 2D case, it is not helpful in identifying the defects responsible for plasticity. In order to solve this problem, we further differentiate the hedgehog topological defects (HTDs) based on their geometric structure as explained in the previous Section. Interestingly, our simulations do not reveal the presence of only ideal radial or hyperbolic defects (Figure 1e-f). Instead, we observe a range of intermediate configurations characterized by different values of $N_{s}$, with representative examples shown in Supplementary Figure 9a-f.
In Figure 3a-c, we present the spatial correlation functions for hyperbolic-hyperbolic, hyperbolic-radial, and radial-radial defect pairs. We observe that hyperbolic-hyperbolic and radial-radial correlations are stronger compared to hyperbolic-radial correlations, indicating a natural tendency for defects of the same geometric type to be spatially correlated. In Figure 3d, we show the correlation functions between soft spots and defects, $g_{s\alpha}(r)$, where $\alpha=H,R$. At low frequencies, the correlation with hyperbolic defects remains consistently stronger than with radial defects, as evident from Figure 3d. To further quantify this behavior, we compute the frequency-weighted average correlation function [69]:
$g_{s\alpha}^{\text{av}}(r)=\frac{\sum_{k}g_{k,s\alpha}(r)/{\omega}_{k}^{2}}{\sum_{k}1/{\omega}_{k}^{2}},$ (8)
where $\alpha=H,R$, and the sum runs over the selected frequency range. The $\omega^{2}$ weighting accounts for the quadratic scaling of the defect density at low frequencies. We consider a frequency range $\omega_{k}\in[1,\omega_{c}]$ with $\omega_{c}=3.5$, discarding $\omega_{k}<1$ due to large fluctuations and the significantly lower number of hedgehog defects ($N_{d}\approx 0-50$) in this range.
The averaged correlation functions $g_{s\alpha}^{\text{av}}(r)$ for $\alpha=H,R$ are shown in Figure 3e, clearly demonstrating that hyperbolic defects are more strongly correlated with soft spots than radial defects. In Figure 3f, we present the averaged correlation functions $g_{s\alpha}^{\text{av}}(r)$ for different values of $N_{s}$ (the number of saddle-like surfaces). The results reveal that the correlation strength increases with $N_{s}$, indicating that defects with more saddle-like features are more strongly associated with soft spots. We further verify this trend by varying the frequency range used in the averaging procedure. As the frequency range $\omega_{c}$ increases, the first peak height in $g_{s\alpha}^{\text{av}}(r)$ decreases, as shown in the supplementary figure 11a. This suggests that the correlation is strongest at lower frequencies, in agreement with previous reports in the literature [45]. Additionally, we compute the averaged correlation function for defects categorized by their topological charge $Q=+1$ and $Q=-1$ and find no significant differences, as shown in the supplementary figure 11b.
These findings highlight that, for 3D point defects in the eigenvector field, geometry, and not only topology, plays a fundamental role in identifying the carriers of plasticity. In particular, our results suggest that the topological defects relevant to plasticity are those with hyperbolic geometry, aligning with the results in 2D systems [45, 72] where this hyperbolic structure is associated to anti-vortices with $q_{2D}=-1$.
Predicting plasticity from the topology of the displacement vector
@@
Inspired by the proposal of looking at topological defects in the dynamical displacement field $\vec{u}$ [44], rather than in the eigenvector field, we now apply the same 3D algorithm to find HTDs in $\vec{u}$. We notice that there is no direct relation between the displacement field and a single eigenvector $\mathbf{e}_{k}$, despite $\vec{u}$ can be decomposed onto the complete basis set of eigenvectors. Therefore, a correlation between topological defects in the eigenvector field and in the displacement field is not obvious. Nevertheless, around plastic events, it is known that the dynamical response is dominated by a few localized and low frequency modes (see e.g. [73]), providing a possible connection between the topology of the low-frequency eigenvectors and that of the displacement field.
@@
In our study, we apply quasi-static shear to our system using the athermal quasi static (AQS) protocol [74] with strain step $\delta\gamma=10^{-3}$. The stress-strain curve is presented in the Methods section (Figure 6c) and displays a yielding instability at $\gamma^{*}\approx 0.099$. For presentation purposes, here we focus on four plastic events, including the yielding point, and consider the non-affine displacement field at these specific strain values. In order to test the universality of our findings, we define the soft spots using two different methods: (I) the softness parameter defined in Eq. (7) (and used in the previous analysis) and (II) the value of the particle level non-affine displacement field $\vec{u}_{\text{NA}}$. Within this second method, we identify the plastic soft spots by looking at the locations of the $5\%$ of particles having higher values of $|\vec{u}_{\text{NA}}|$.
@@
Figure 4: Correlation between soft spots and HTDs. a Spatial correlation functions between soft spots (identified using the softness parameter, Eq. (7)) and HTDs within the non-affine displacement field. The correlations for HTDs with topological charge $Q=+1$ ($-1$) are shown in red (blue), respectively. b Spatial correlations between soft spots and HTDs with hyperbolic and radial geometrical nature. Both data sets are averaged over different $\gamma$ values (with error bars), corresponding to plastic events as indicated in Figure 6c.